# Prospects for Cosmological Research with the FAST Array: 21-cm Intensity Mapping Survey Observation Strategies

### v1

@@ Abstract
Precise cosmological measurements are essential for understanding the evolution of the universe and the nature of dark energy. The Five-hundred-meter Aperture Spherical Telescope (FAST), the most sensitive single-dish radio telescope, has the potential to provide the precise cosmological measurements through neutral hydrogen 21-cm intensity mapping sky survey. This paper primarily explores the potential of technological upgrades for FAST in cosmology. The most crucial upgrade begins with equipping FAST with a wide-band receiver ($2.50<z<$ 2.5). This upgrade can enable FAST to achieve higher precision in cosmological parameter estimation than the Square Kilometre Array Phase-1 Mid frequency. On this basis, expanding to a FAST array (FASTA) consisting of six identical FAST would offer significant improvements in precision compared to FAST. Additionally, compared with the current results from the data combination of cosmic microwave background, baryon acoustic oscillations (optical galaxy surveys), and type Ia supernovae, FASTA can provide comparable constraints. Specifically, for the dark-energy equation-of-state parameters, FASTA can achieve $\sigma(w_{0})=0.09$ and $\sigma(w_{a})=0.33$.
I Introduction
@@
Dark energy is a component of the universe that causes the late-time cosmic acceleration. Although it is known to comprise 68% of the total energy density of the universe, the nature of dark energy remains unknown. The $\Lambda$ cold dark matter ($\Lambda$CDM) model, the standard model of cosmology, treats dark energy as a constant energy density filling space homogeneously. However, the latest baryon acoustic oscillation (BAO) observation from the Dark Energy Spectroscopic Instrument (DESI) suggests that dark energy may potentially be a dynamical energy density evolving with time Aghamousa et al. (2016); Adame et al. (2024a, b). This finding indicates the possibility of dark-energy dynamical evolution, but more observations and advanced instruments are needed to fully confirm this and to understand the evolutionary dynamics of dark energy.
@@
BAO from the large-scale structure (LSS) is a late-time cosmological probe that can be used as a standard ruler to study the expansion history of the universe Blake and Glazebrook (2003); Cole et al. (2005); Percival et al. (2010); Beutler et al. (2011); Blake et al. (2011); Ross et al. (2015); Alam et al. (2017); Ata et al. (2018). Currently, BAO studies primarily rely on optical galaxy surveys York et al. (2000); Percival et al. (2001); Kaiser et al. (2002); Collaboration: et al. (2016); Aghamousa et al. (2016). While effective, these surveys are limited in terms of efficiency and coverage. To overcome these limitations and achieve more precise cosmological measurements, advancements in radio astronomy have developed a promising method known as 21 cm intensity mapping (21 cm IM). After reionization, neutral hydrogen ($\mathrm{H\ {\scriptstyle I}}$) mainly exists within galaxies. This technique measures the intensity of the characteristic 21 cm emission line from $\mathrm{H\ {\scriptstyle I}}$ to study the LSS Furlanetto et al. (2006); Chang et al. (2010); Switzer et al. (2013); Masui et al. (2013); Pritchard and Loeb (2012). This method does not require resolving individual galaxies, thus it has high survey efficiency. The characteristics of the 21 cm signal allow us to probe deeper into the universe and directly obtain redshift information. Recent forecasts and observational studies have confirmed the potential of 21 cm IM in cosmology Bull et al. (2015); Zhang et al. (2019, 2020); Jin et al. (2020); Hu et al. (2020); Zhang et al. (2021); Jin et al. (2021); Wu and Zhang (2022); Wu et al. (2023a, b); Chen (2023); Zhang et al. (2023a); Li et al. (2023); Cunnington et al. (2022).
@@
To achieve more precise cosmological measurements, advanced radio telescopes are utilized in 21 cm IM. These telescopes include those currently in use, such as the Five-hundred-meter Aperture Spherical Telescope (FAST) Nan et al. (2011), the MeerKAT Santos et al. (2016); Wang et al. (2021); Li et al. (2020a), the Tianlai Chen (2011, 2012); Li et al. (2020b); Wu et al. (2021); Perdereau et al. (2022); Sun et al. (2022), the Canadian Hydrogen Intensity Mapping Experiment (CHIME) Newburgh et al. (2014), and those planned for future use, such as the Square Kilometre Array (SKA) Dewdney et al. (2009); Santos et al. (2015); Braun et al. (2015); Bacon et al. (2020); An et al. (2022), the Baryon Acoustic Oscillations from Integrated Neutral Gas Observations (BINGO) Battye et al. (2012); Wuensche (2019), and the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX) Newburgh et al. (2016).
@@
Among these, FAST is particularly significant due to its large aperture and sensitivity. The current FAST has achieved substantial success in many areas, particularly in fast radio burst and gravitational wave research Li et al. (2021); Niu et al. (2022); Xu et al. (2023). However, to harness its potential for 21 cm IM and to extend its capability for cosmological studies, several upgrades are planned for FAST Zhang et al. (2023b); Xue et al. (2023); Yin et al. (2023).
@@
This paper aims to explore how potential future upgrades to FAST could promote cosmological research by using 21 cm IM. For the expansion of FAST itself, equipping it with a wide-band receiver aims to extend its 21 cm signal detection capability, enabling broader redshift coverage Zhang et al. (2023b). This will allow us to observe a larger volume of the universe, aiding in constructing a more precise three-dimensional map of the LSS. Moreover, a phased array feed (PAF) is planned to enhance its survey efficiency Landon et al. (2010); Hotan et al. (2021). Additionally, the expansion of the construction around FAST is also considered. The plan involves building 40-meter diameter spherical radio telescopes (40 m dishes) in two phases around FAST: the first phase involves constructing six 40 m dishes, and the second phase will increase this number to twenty. The most significant upgrade is the FAST Array (FASTA), which plans to add five more identical 500-meter diameter telescopes to join the current FAST Xue et al. (2023). This network will significantly improve sensitivity and survey efficiency, enhancing the ability to explore LSS and providing higher precision in cosmological parameter estimation.
@@
In Section II, we introduce the 21 cm IM simulation and the methods used to constrain cosmological parameters. In Section III, we discuss the performance of various configurations for 21 cm IM and the resulting cosmological constraints. In Section IV, we conclude with perspectives for future research.
@@
In our analysis, we adopt the Planck best-fit $\Lambda$CDM model as fiducial cosmology Collaboration (2020). The parameters are $H_{0}=67.3$ km s-1 Mpc-1, $\Omega_{\mathrm{m}}=0.317$, $\Omega_{\mathrm{b}}=0.0495$, $\Omega_{k}=0$, $\sigma_{8}=0.812$, and $n_{\mathrm{s}}=0.965$.
II 21 cm INTENSITY MAPPING
II.1 Signal
@@
21 cm IM measures the intensity of 21 cm signals and converts it into brightness temperature. The mean temperature of $\mathrm{H\ {\scriptstyle I}}$ can be expressed as Bull et al. (2015)
$\overline{T}_{\mathrm{b}}=\frac{3}{32\pi}\frac{\mathcal{h}c^{3}A_{10}}{k_{\mathrm{B}}m_{\mathrm{p}}\nu_{21}^{2}}\frac{(1+z)^{2}}{H(z)}\Omega_{\mathrm{H\,{\scriptscriptstyle I}}}(z)\rho_{\mathrm{c},0}\ ,$ (1)
@@
where $\mathcal{h}$ is the Planck constant, $c$ is the speed of light, $A_{10}$ is the Einstein coefficient for spontaneous emission Furlanetto et al. (2006), $k_{\mathrm{B}}$ is the Boltzmann constant, $m_{\mathrm{p}}$ is the proton mass, $\nu_{21}$ is the $\mathrm{H\ {\scriptstyle I}}$ line frequency, $H(z)$ is the Hubble parameter, $\Omega_{\mathrm{H\,{\scriptscriptstyle I}}}(z)$ is the fractional density of $\mathrm{H\ {\scriptstyle I}}$, and $\rho_{\mathrm{c},0}$ is the current critical density of the universe.
The signal covariance is expressed in terms of the power spectrum Blake and Glazebrook (2003); Bull et al. (2015)
$C^{\mathrm{S}}(\bm{q},y)=\frac{T_{b}^{2}(z_{i})}{r^{2}r_{\nu}}(b_{\mathrm{H\,{\scriptscriptstyle I}}}+f\mu^{2})^{2}\exp(-k^{2}\mu^{2}\sigma_{\mathrm{NL}}^{2})P(k,z),$ (2)
@@
where $\bm{q}=\bm{k}_{\perp}r$ and $y=k_{\|}r_{\nu}$. $r$ denotes the comoving distance to $z$, and $r_{\nu}$ corresponds to the comoving distance for the frequency interval. $b_{\mathrm{H\,{\scriptscriptstyle I}}}$ represents the bias factor of $\mathrm{H\ {\scriptstyle I}}$ Xu et al. (2014). $f$ is the linear growth rate of the structure. $\mu=k_{\|}/k$ is the cosine of the angle between the wavevector $\bm{k}$ and the line of sight. $\sigma_{\mathrm{NL}}=7\ \mathrm{Mpc}$ represents the non-linear dispersion scale Li et al. (2007). $P(k,z)$ is the matter power spectrum at wavenumber $k$ and $z$, which can be calculated by CAMB Lewis et al. (2000).
II.2 Noise
The noise covariance can be written as Bull et al. (2015)
$C^{\mathrm{N}}(\bm{q},y)=\frac{T_{\text{sys }}^{2}}{t_{\text{tot }}\Delta\nu\ N_{\mathrm{b}}N_{\mathrm{d}}}U_{\text{bin }}B_{\perp}^{-2}B_{\|}^{-1},$ (3)
@@ II.2 Noise
where $T_{\mathrm{sys}}$ includes the instrumental temperature $T_{\mathrm{inst}}$ and the sky background radiation $T_{\text{sky}}\approx 60\mathrm{~{}K}\times(\nu/300\ \mathrm{MHz})^{-2.5}$. $t_{\mathrm{tot}}$ represents the total integration time, $N_{\mathrm{b}}$ is the number of beams, $N_{\mathrm{d}}$ is the number of dishes in the array, and $\Delta\nu$ is the total bandwidth. $U_{\text{bin}}=S_{\text{area}}\Delta\tilde{\nu}$ represents the survey volume element at a given redshift, where $S_{\text{area}}$ is the survey area and $\Delta\tilde{\nu}=\Delta{\nu}/\nu_{21}$ is the dimensionless bandwidth for a given redshift bin. $B_{\|}$ is the radial beam response
$B_{\|}(y)=\exp\left(-\frac{\left(y\delta\nu/\nu_{21}\right)^{2}}{16\ln 2}\right),$ (4)
@@
$B_{\perp}(\bm{q})=\exp\left(-\frac{\left(q\frac{\lambda}{D_{\mathrm{dish}}}\right)^{2}}{16\ln 2}\right),$ (5)
@@
II.3 Foreground
The foreground is an important component of the signal received by the telescope. It primarily originates from the galaxy and extragalactic point sources. At this stage, the effectiveness of methods for subtracting foregrounds from observational data has been very limited. The mainstream methods mainly include Polynomial Fitting, Principal Component Analysis, Independent Component Analysis, and Artificial Intelligence Wang et al. (2006); Morales et al. (2012); Parsons et al. (2012); Liu et al. (2014); Shaw et al. (2015); Zhu et al. (2018); Zuo et al. (2018); Carucci et al. (2020); Cunnington et al. (2021); Ni et al. (2022); Gao et al. (2023). We assume a method that can effectively subtract the foreground, and the residual foreground model can be written as Bull et al. (2015); Santos et al. (2005)
$C^{\mathrm{F}}(\bm{q},y)=\varepsilon_{\mathrm{FG}}^{2}\sum_{X}A_{X}\left(\frac{l_{p}}{2\pi q}\right)^{n_{X}}\left(\frac{\nu_{p}}{\nu_{i}}\right)^{m_{X}},$ (6)
where $\varepsilon_{\mathrm{FG}}$ is a factor that describes the efficiency of foreground subtraction. $\varepsilon_{\mathrm{FG}}=1$ represents no foreground subtraction, while $\varepsilon_{\mathrm{FG}}=0$ represents perfect foreground subtraction. We assume $\varepsilon_{\mathrm{FG}}=10^{-6}$, as this level is effective for extracting cosmological signals. $X$ is the different foreground component shown in Table 1.
Table 1: Foreground model parameters at $l_{p}=1000$ and $\nu_{p}=130$ MHz Santos et al. (2005).

### v2

@@ Abstract
Precise cosmological measurements are essential for understanding the evolution of the universe and the nature of dark energy. The Five-hundred-meter Aperture Spherical Telescope (FAST), the most sensitive single-dish radio telescope, has the potential to provide the precise cosmological measurements through neutral hydrogen 21 cm intensity mapping sky survey. This paper primarily explores the potential of technological upgrades for FAST in cosmology. The most crucial upgrade begins with equipping FAST with a wide-band receiver ($2.50<z<$ 2.5). This upgrade can enable FAST to achieve higher precision in cosmological parameter estimation than the Square Kilometre Array Phase 1 Mid-Frequency Array. On this basis, expanding to a FAST array (FASTA) consisting of six identical FASTs would offer significant improvements in precision compared to FAST. Additionally, compared with the current results from the data combination of cosmic microwave background, baryon acoustic oscillations (optical galaxy surveys), and type Ia supernovae, FASTA can provide comparable constraints. Specifically, for the dark-energy equation-of-state parameters, FASTA can achieve $\sigma(w_{0})=0.09$ and $\sigma(w_{a})=0.33$.
1 Introduction
@@
Dark energy is a component of the universe that causes the late-time cosmic acceleration. Although it is known to comprise 68% of the total energy density of the universe, the nature of dark energy remains unknown. The $\Lambda$ cold dark matter ($\Lambda$CDM) model, the standard model of cosmology, treats dark energy as a constant energy density filling space homogeneously. However, recent observations of baryon acoustic oscillations (BAO) from the Dark Energy Spectroscopic Instrument (DESI) have shown deviations from the $\Lambda$CDM model [1, 2, 3, 4, 5, 6, 7, 8]. The result suggests that dark energy may be a dynamical energy density evolving over time or indicate the presence of other new physics. Nonetheless, further observations and advanced instruments are needed to confirm these possibilities and to better understand the evolutionary dynamics of dark energy.
@@
BAO from the large-scale structure (LSS) is a late-time cosmological probe that can be used as a standard ruler to study the expansion history of the universe [9, 10, 11, 12, 13, 14, 15, 16]. Currently, BAO studies primarily rely on optical galaxy surveys [17, 18, 19, 20, 1]. While effective, these surveys are limited in terms of efficiency and coverage. To overcome these limitations and achieve more precise cosmological measurements, advancements in radio astronomy have developed a promising method known as 21 cm intensity mapping (21 cm IM). After reionization, neutral hydrogen ($\mathrm{H\ {\scriptstyle I}}$) mainly exists within galaxies. This technique measures the intensity of the characteristic 21 cm emission line from $\mathrm{H\ {\scriptstyle I}}$ to study the LSS [21, 22, 23, 24, 25]. This method does not require resolving individual galaxies, thus it has high survey efficiency. The characteristics of the 21 cm signal allow us to probe deeper into the universe and directly obtain redshift information. Recent forecasts and observational studies have confirmed the potential of 21 cm IM in cosmology [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. In addition to the 21 cm line, intensity mapping of other spectral lines, such as CO and CII, has also attracted attention. These lines can provide information about the LSS and the distribution of molecular and ionized gas in galaxies [40, 41, 42].
@@
To achieve more precise cosmological measurements, advanced radio telescopes are utilized in 21 cm IM. These telescopes include those currently in use, such as the Five-hundred-meter Aperture Spherical Telescope (FAST) [43], the MeerKAT [44, 45, 46], the Tianlai [47, 48, 49, 50, 51, 52], the Canadian Hydrogen Intensity Mapping Experiment (CHIME) [53]. Currently, 21 cm signals have been observed through cross-correlation with galaxy surveys [22, 24, 54, 46, 55, 56, 57], and the MeerKAT telescope has also detected the autocorrelation spectrum on Mpc scales [58]. However, the large scale autocorrelation power spectrum still has not been observed [23]. Future instruments planned for use include the Square Kilometre Array (SKA) [59, 60, 61, 62, 63], the Baryon Acoustic Oscillations from Integrated Neutral Gas Observations (BINGO) [64, 65], the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX) [66], and the Canadian Hydrogen Observatory and Radio-transient Detector (CHORD) [67]. These advanced facilities are expected to further enhance the capabilities of 21 cm IM, enabling more detailed and accurate cosmological studies.
@@
Among these, FAST is particularly significant due to its large aperture and sensitivity. The current FAST has achieved substantial success in many areas, particularly in fast radio burst and gravitational wave researches [68, 69, 70]. However, to harness its potential for 21 cm IM and to extend its capability for cosmological studies, several upgrades are planned for FAST [71, 72, 73, 74].
@@
This paper aims to explore how potential future upgrades to FAST could promote cosmological research by using 21 cm IM. For the expansion of FAST itself, equipping it with a wide-band receiver aims to extend its 21 cm signal detection capability, enabling broader redshift coverage [71]. This will allow us to observe a larger volume of the universe, aiding in constructing a more precise three-dimensional map of the LSS. Moreover, a phased array feed (PAF) is planned to enhance its survey efficiency [75, 76]. Additionally, the expansion of the construction around FAST is also considered. The plan involves building 40-meter aperture radio telescopes (40-m antennas) in two phases around FAST: the first phase involves constructing six 40-m antennas, and the second phase will increase this number to twenty-four [74]. The most significant upgrade is the FAST array (FASTA), which plans to add five more identical 500-meter diameter telescopes to join the current FAST [72]. This network will significantly improve sensitivity and survey efficiency, enhancing the ability to explore LSS and providing higher precision in cosmological parameter estimation.
@@
In Section 2, we introduce the 21 cm IM simulation and the methods used to constrain cosmological parameters. In Section 3, we discuss the performance of various configurations for 21 cm IM and the resulting cosmological constraints. In Section 4, we conclude with perspectives for future research.
@@
In our analysis, we adopt the Planck best-fit $\Lambda$CDM model as fiducial cosmology [77]. The parameters are $H_{0}=67.3$ km s-1 Mpc-1, $\Omega_{\mathrm{m}}=0.317$, $\Omega_{\mathrm{b}}=0.0495$, $\Omega_{K}=0$, $\sigma_{8}=0.812$, and $n_{\mathrm{s}}=0.965$.
2 21 cm intensity mapping
2.1 Signal
@@
21 cm IM measures the intensity of 21 cm signals and converts it into brightness temperature. The mean temperature of $\mathrm{H\ {\scriptstyle I}}$ can be expressed as [26]
$\overline{T}_{\mathrm{b}}=\frac{3}{32\pi}\frac{\mathcal{h}c^{3}A_{10}}{k_{\mathrm{B}}m_{\mathrm{p}}\nu_{21}^{2}}\frac{(1+z)^{2}}{H(z)}\Omega_{\mathrm{H\,{\scriptscriptstyle I}}}(z)\rho_{\mathrm{c},0}\ ,$ (2.1)
@@
where $\mathcal{h}$ is the Planck constant, $c$ is the speed of light, $A_{10}$ is the Einstein coefficient for spontaneous emission [21], $k_{\mathrm{B}}$ is the Boltzmann constant, $m_{\mathrm{p}}$ is the proton mass, $\nu_{21}$ is the $\mathrm{H\ {\scriptstyle I}}$ line frequency, $H(z)$ is the Hubble parameter, $\Omega_{\mathrm{H\,{\scriptscriptstyle I}}}(z)$ is the fractional density of $\mathrm{H\ {\scriptstyle I}}$, and $\rho_{\mathrm{c},0}$ is the current critical density of the universe.
The signal covariance is expressed in terms of the power spectrum [9, 26]
$C^{\mathrm{S}}(\bm{q},y)=\frac{T_{b}^{2}(z_{i})}{r^{2}r_{\nu}}(b_{\mathrm{H\,{\scriptscriptstyle I}}}+f\mu^{2})^{2}\exp(-k^{2}\mu^{2}\sigma_{\mathrm{NL}}^{2})P(k,z),$ (2.2)
@@
where $\bm{q}=\bm{k}_{\perp}r$ and $y=k_{\|}r_{\nu}$. $r$ denotes the comoving distance to $z$, and $r_{\nu}$ corresponds to the comoving distance for the frequency interval. $b_{\mathrm{H\,{\scriptscriptstyle I}}}$ represents the bias factor of $\mathrm{H\ {\scriptstyle I}}$ [78]. $f$ is the linear growth rate of the structure. $\mu=k_{\|}/k$ is the cosine of the angle between the wavevector $\bm{k}$ and the line of sight. $\sigma_{\mathrm{NL}}=7\ \mathrm{Mpc}$ represents the non-linear dispersion scale [79]. $P(k,z)$ is the matter power spectrum at wavenumber $k$ and redshift $z$, which can be calculated by CAMB [80].
2.2 Noise
The noise covariance can be written as [26]
$C^{\mathrm{N}}(\bm{q},y)=\frac{T_{\text{sys }}^{2}}{t_{\text{tot }}\Delta\nu\ N_{\mathrm{b}}N_{\mathrm{d}}}U_{\text{bin }}B_{\perp}^{-2}B_{\|}^{-1},$ (2.3)
@@ 2.2 Noise
where $T_{\mathrm{sys}}$ includes the instrumental temperature $T_{\mathrm{inst}}$ and the sky background radiation $T_{\text{sky}}\approx 60\mathrm{~{}K}\times(\nu/300\ \mathrm{MHz})^{-2.5}$. $t_{\mathrm{tot}}$ represents the total integration time, $N_{\mathrm{b}}$ is the number of beams, $N_{\mathrm{d}}$ is the number of dishes in the array, and $\Delta\nu$ is the total bandwidth. $U_{\text{bin}}=S_{\text{area}}\Delta\tilde{\nu}$ represents the survey volume at a given redshift, where $S_{\text{area}}$ is the survey area and $\Delta\tilde{\nu}=\Delta{\nu}/\nu_{21}$ is the dimensionless bandwidth for a given redshift bin. $B_{\|}$ is the radial beam response
$B_{\|}(y)=\exp\left(-\frac{\left(y\delta\nu/\nu_{21}\right)^{2}}{16\ln 2}\right),$ (2.4)
@@
$B_{\perp}(\bm{q})=\exp\left(-\frac{\left(q\frac{\lambda}{D_{\mathrm{dish}}}\right)^{2}}{16\ln 2}\right),$ (2.5)
@@
2.3 Foreground
The foreground is an important component of the signal received by the telescope. It primarily originates from our own galaxy and extragalactic point sources. At this stage, the effectiveness of methods for subtracting foregrounds from observational data has been very limited. The methods include Polynomial Fitting, Principal Component Analysis, Independent Component Analysis, and others [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. We assume a method that can effectively subtract the foreground, and the residual foreground model can be written as [26, 93]
$C^{\mathrm{F}}(\bm{q},y)=\varepsilon_{\mathrm{FG}}^{2}\sum_{X}A_{X}\left(\frac{l_{p}}{2\pi q}\right)^{n_{X}}\left(\frac{\nu_{p}}{\nu_{i}}\right)^{m_{X}},$ (2.6)
where $\varepsilon_{\mathrm{FG}}$ is a factor that describes the efficiency of foreground subtraction. $\varepsilon_{\mathrm{FG}}=1$ represents no foreground subtraction, while $\varepsilon_{\mathrm{FG}}=0$ represents perfect foreground subtraction. We assume $\varepsilon_{\mathrm{FG}}=10^{-6}$, as this level is effective for extracting cosmological signals. $X$ is the varies foreground component shown in Table 1.
Table 1: Foreground model parameters at $l_{p}=1000$ and $\nu_{p}=130$ MHz [93].
@@ 2.3 Foreground
2.4 Instrument parameters
In this paper, we consider several different configurations of FAST and the Square Kilometre Array Phase 1 Mid-Frequency Array (SKA1-Mid) for simulating the 21 cm IM. The specific instrument parameters are listed in Table 2.
The current capability of FAST to detect 21 cm signals covers the redshift range $0.350<z<$ 0.35. A new wide-band receiver under test extends this range to $1.840<z<$ 1.84 [71]. Currently, DESI measurements have mapped the cosmological distance-redshift relation with an accuracy of 1%–3% across redshifts below $z<2.5$ [2]. This paper assumes that the upgraded new wide-band receiver on FAST can detect 21 cm signals up to $2.50<z<$ 2.5.
Additionally, constructing 6$\times$40-m antennas or 24$\times$40-m antennas around FAST is considered. These dishes will use the same receiver and feed system as FAST, so their frequency bands and $N_{\mathrm{b}}$ will be identical. Given the outstanding performance of FAST, we also consider constructing five additional identical FAST telescopes to form FASTA, significantly improving survey efficiency. Furthermore, we consider equipping FAST with a PAF to increase the value of $N_{\mathrm{b}}$ to 100.
Regarding instrument temperature, in Ref. [94] the authors measured the temperature of FAST near boresight to be 20 K in the L-band. We assume that the frequency band can be broadened while maintaining the instrument temperature at 20 K. For the PAF, we assume the temperature to be approximately 30 K due to the increase in the number of beams. The total observation time is assumed to be 10000 hours. The calculation for the survey area is as follows
$S_{\text{area}}=\eta\int_{\phi}\int_{\theta}\sin\theta\mathrm{d}\theta\mathrm{d}\phi\ ,$ (2.7)
where $\eta$ represents the fraction of the initial survey area that remains usable for scientific observation after excluding the Milky Way. For FAST/FASTA, we consider $\eta=0.7$ and the range of declination from $-14^{\circ}$ to $+66^{\circ}$, yielding an area of approximately 8100 $\text{deg}^{2}$.
For SKA1-Mid, we consider Band 1 ($30.35<z<$ 3) planned for 21 cm IM, with the survey area of 20000 deg2 and the total observation time of 10000 hours. The instrument temperature is modeled based on actual measurements and simulation data from the SKA Observatory, and the form is as follows [62, 95]
$T_{\mathrm{inst}}=15\mathrm{~{}K}+30\mathrm{~{}K}\left(\frac{f}{\mathrm{GHz}}-0.75\right)^{2}.$ (2.8)